# Diffusion Model as a Noise-Aware Latent Reward Model for Step-Level Preference Optimization

### v2

@@ Appendix B Additional Experiments
Regularization HyperParameter $\beta$. $\beta$ is a regularization hyperparameter to control the deviation of the optimized model ($p_{\theta}$ in Eqn. (7)) with the reference model $p_{ref}$. As shown in Tab. 13, as $\beta$ increases, the regularization effect becomes stronger, which slows down the model’s optimization speed and leads to poorer performance. Conversely, if $\beta$ is too small ($\beta=20$), the regularization constraint becomes too weak, potentially reducing the model’s generalization ability and decreasing the quality of the generated images. The performance is optimal when $\beta$ is within the range of 100 to 500.
Table 14: Ablataion results on the dynamic threshold strategies.
@@
Other Dynamic Threshold Strategy. In the main paper, we dynamically correlate the threshold in LPO sampling with the standard deviation $\sigma_{t}$ in DDPM (Ho et al., 2020). Here we explore alternative strategies. The first strategy involves replacing $\sigma_{t}$ with variance $\sigma_{t}^{2}$. Consequently, the threshold can be formulated as:
$th_{t}=\frac{t-t_{min}}{t_{max}-t_{min}}*(th_{max}-th_{min})+th_{min},$ (15)
Figure 7: The thresholds of different strategies.
Figure 8: The number of valid samples of different timesteps .
Fig. 8 illustrates the number of valid samples of different timesteps for both constant and dynamic (standard deviation) threshold strategies. Dynamic threshold strategy significantly increases the number of valid samples, especially at intermediate timesteps.
Appendix C Additional Visualizations
Predicted Images $I_{t}$ at Different Timesteps. Fig. 9 illustrates the predicted images $I_{t}$ generated through the diffusion forward and VAE decoding processes, as shown in Fig. 3 (a). As discussed in the main paper, predicted images at large timesteps tend to be very blurred, causing a significant distribution shift from the original images. This makes it challenging for PRMs to adapt to these blurred images without adequate pre-training and extensive datasets, resulting in unreliable predictions at large timesteps.
@@ Appendix C Additional Visualizations
More Comparison. Fig. 11 presents the generated images of different methods based on SD1.5 (Rombach et al., 2022). Fig. 12 and Fig. 13 show the larger version of images in Fig. 2. Fig. 14 and Fig. 15 provide more generated images of various methods based on SDXL (Podell et al., 2023). Some keywords in prompts are emphasized using bold formatting.
Figure 9: Predicted images $I_{t}$ in Fig. 3 (a) at different timesteps. The original images come from the Pick-a-Pic v1 dataset (Kirstain et al., 2023).

### v3

@@ Appendix C Additional Ablation Experiments
The Compatibility with High Noise at Large Timesteps. In Tab. 23, we analyze the compatibility of SPO and LPO with large training timesteps and high noises. Notably, for SPO, the training efficiency within the timestep range [0, 950] is significantly lower than that at [0, 750], and it is even inferior to the original model. In contrast, LPO optimization with this range remarkably enhances the model’s performance. Furthermore, training only with the range [750, 950] can even achieve comparable performance to that of the range [0, 950], illustrating the superior adaptability of LRM to high-intensity noise at large timesteps.
LRM’s Sensitivity to Timesteps. In Tab. 23, we investigate the timestep sensitivity of LRM. When LRM is consistently provided with a fixed timestep of 0 during LPO sampling, the model’s performance degrades significantly. This highlights the importance of timestep input and confirms that LRM effectively leverages timestep information to predict preferences for noisy latent images.
Regularization HyperParameter $\beta$. $\beta$ is a regularization hyperparameter to control the deviation of the optimized model ($p_{\theta}$ in Eqn. (6)) with the reference model $p_{ref}$. As shown in Tab. 23, as $\beta$ increases, the regularization effect becomes stronger, which slows down the model’s optimization speed and leads to poorer performance. Conversely, if $\beta$ is too small ($\beta=20$), the regularization constraint becomes too weak, potentially harming the model’s generalization and reducing image quality. The performance is optimal when $\beta$ is within the range of 100 to 500.
Figure 7: The thresholds of different strategies.
Figure 8: The number of valid samples of different timesteps .
Other Dynamic Threshold Strategy. In the main paper, we dynamically correlate the threshold in LPO sampling with the standard deviation $\sigma_{t}$ in DDPM [44]. Here we explore alternative strategies. The first strategy involves replacing $\sigma_{t}$ with variance $\sigma_{t}^{2}$. Consequently, the threshold can be formulated as:
$th_{t}=\frac{t-t_{min}}{t_{max}-t_{min}}*(th_{max}-th_{min})+th_{min},$ (17)
Fig. 8 illustrates the number of valid samples of different timesteps for both constant and dynamic (standard deviation) threshold strategies. The dynamic threshold strategy significantly increases the number of valid samples, especially at intermediate timesteps.
Complete Results on GenEval. The complete GenEval results corresponding with that in Tab. 2 are presented in Tab. 26 and Tab. 26.
Table 24: Ablation results on the dynamic threshold strategies.
@@
Table 25: Complete results on GenEval [40] with 20 inference steps.
SD1.5 [14] Original 97.50 37.12 34.69 75.53 3.75 6.75 42.56
Diff.-DPO [4] 98.44 38.38 36.25 77.93 4.50 7.25 43.79
SPO [8] 95.00 33.84 32.50 69.95 4.25 7.25 40.46
SePPO [30] 99.06 39.90 34.69 81.91 6.25 8.00 44.97
LPO 97.81 54.80 40.94 79.52 7.00 10.25 48.39
SDXL [11] Original 93.75 63.38 30.94 80.05 9.25 19.00 49.40
Diff.-DPO [4] 99.06 76.52 45.00 88.83 11.50 25.75 57.78
MaPO [29] 95.63 68.94 32.19 83.51 11.50 17.75 51.59
SPO [8] 94.38 69.44 31.88 81.65 10.25 15.50 50.52
InterComp [31] 99.06 82.83 43.75 84.04 14.00 31.75 59.24
LPO 99.69 81.57 43.75 89.10 14.00 27.50 59.27
Table 26: Complete results on GenEval [40] with 50 inference steps.
SD1.5 [14] Original 98.13 37.88 33.13 76.33 3.75 6.00 42.53
Diff.-DPO [4] 98.13 41.16 37.81 81.91 4.50 7.25 45.13
SPO [8] 95.63 36.62 34.83 72.34 3.75 6.50 41.53
SePPO [30] 98.13 41.16 33.44 79.52 7.25 7.25 44.46
LPO 97.81 55.30 42.19 80.59 6.75 10.00 48.77
SDXL [11] Original 94.38 67.68 41.56 81.65 10.50 18.00 52.29
Diff.-DPO [4] 99.06 77.78 49.69 86.17 13.25 27.50 58.91
MaPO [29] 96.56 66.41 40.00 84.31 10.75 18.75 52.80
SPO 97.81 73.48 41.25 85.64 13.00 20.00 55.20
SPO [8] 96.88 69.70 37.19 83.51 9.50 19.75 52.75
InterComp [31] 99.06 85.10 41.25 86.97 13.50 32.00 59.64
LPO 99.69 84.34 43.13 90.43 13.75 27.75 59.85
Appendix D Additional Visualizations
Predicted Images $I_{t}$ at Different Timesteps. Fig. 9 illustrates the predicted images $I_{t}$ generated through the diffusion forward and VAE decoding processes, as shown in Fig. 4 (a). As discussed in the main paper, predicted images at large timesteps tend to be very blurred, causing a significant distribution shift from the original images. This makes it challenging for PRMs to adapt to these blurred images without adequate pre-training and extensive datasets, resulting in unreliable predictions at large timesteps. In contrast, LRM can naturally perceive noisy latent images even at large timesteps, as it does during pre-training, representing a significant advantage over PRMs.
@@ Appendix D Additional Visualizations
More Comparison. Fig. 11 presents the generated images of different methods based on SD1.5 [14]. Fig. 12 and Fig. 13 show the larger version of images in Fig. 1 with corresponding prompts. Fig. 14 and Fig. 15 provide more generated images of various methods based on SDXL [11]. Some keywords in prompts that other models fail to depict are highlighted using bold formatting.
Figure 9: Predicted images $I_{t}$ in Fig. 4 (a) at different timesteps. The original images come from the Pick-a-Pic v1 dataset [16].
@@
Appendix E Broader Impacts
This work introduces a preference optimization method for text-to-image diffusion models. The method may have the following impacts: